OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17

OpenStudy (anonymous):

If the line is perependicular then the product of their slopes is equal to -1 m1m2=-1 Here the slope of the line is y=-3x+7 m1=-3 m2=1/3 There fore the required eqn is y-6=1/3(x+1) 3y-18=x+1 x-3y+19=0